The sum of the ages of Mike and Dan is 20 years. In four years, Dan’s age will be 3/4 of Mike’s age. How old is Mike now?
Answer by Alan3354(69443)

The sum of the ages of Mike and Dan is 20 years. In four years, Dan’s age will be 3/4 of Mike’s age. How old is Mike now?
----------------
M + D = 20
D + 4 = (3/4)*(M + 4)
----
can you do the rest?